OpenStudy (anonymous):

Write a system of equations to solve the following problem. Let c be the number of child tickets and a be the number of adult tickets. Each child ticket for a ride costs $3, while each adult ticket costs $5. If the ride collected a total of $115, and 33 tickets were sold, how many of each type of ticket were sold? A. 3c + 5a = 115 a + c = 33 B. 5c + 3a = 115 c + a = 33 C. 3c + 5a = 33 a + c = 115 D. 3a + 5c = 33 a + c = 115

OpenStudy (anonymous):

i would go with the first one

OpenStudy (anonymous):

others are silly

OpenStudy (anonymous):

if each child's ticket is $3 and there are 4 child tickets sold, then the total is \(3\times 4\) if each child ticket is $3 and there are "c" child tickets sold, the total is \(3c\) for the childs tickets

OpenStudy (anonymous):

similarly \(5a\) for the adult tickets, for a total of \(3c+5a\) which you know is $115

OpenStudy (anonymous):

let "a" be the amount of adult tickets sold let "c" be teh amount of children tickets sold if the total amount of tickets sold is 33 that would mean a+c= 33 now a adult ticket cost $5 and a children ticket costs $3 so the total cost of all the children tickets sold would be 3c whereas the total cost of the adult tickets sold would be 5a if a total of $115 was collected, that would mean 5a+3c=115 i am disappointed at you satellite

OpenStudy (anonymous):

oh wait nvm you did explain it